OpenStudy (hkmiu):

Can someone explain to me how to find all solutions for a triangle with a = 14, b = 12, A = 90 degrees

OpenStudy (hkmiu):

I think there is no solution..

OpenStudy (radar):

It is a right triangle, and you are given two sides, there must be a solution.

OpenStudy (hkmiu):

whoops i totally disregarded that.

OpenStudy (radar):



OpenStudy (hkmiu):

thank you.

OpenStudy (radar):

I will standby while you solve, do you need to find the value of the 3rd side and the remaining 2 angles?

OpenStudy (hkmiu):

so in this case, I can use the Pythagorean Theorem to solve for the sides

OpenStudy (radar):

Yes, and the angle functions for the angles. The 3rd side is irrational \[2\sqrt{13}\]

OpenStudy (radar):



OpenStudy (hkmiu):

it doesn't seem to match up :O

OpenStudy (radar):

Because you stated that angle A was 90 degrees

OpenStudy (radar):

Let me re do it according to the link.

OpenStudy (hkmiu):

sin90/12 = sinA/14

OpenStudy (hkmiu):

I think im just confusing myself.